On the menu list of a restaurant, there are 4 vegetables, 3 salads and 2 types of bread. If a customer wants one vegetable, one salad and one bread, how many possible choices are there for him?


1
Expert's answer
2022-03-19T02:38:22-0400

"N=\begin{pmatrix} 4\\ 1\end{pmatrix}\cdot\begin{pmatrix} 3\\ 1\end{pmatrix}\cdot\begin{pmatrix} 2\\ 1\end{pmatrix}=\\=4\cdot3\cdot2=24."